OpenStudy (anonymous):

Which expressions would complete this equation so that the equation has no solution? 6(4x – 3) – 10x = Choose exactly two answers that are correct. A. 14x – 15 B. 2(7x – 10) C. 15x – 14 D. 2(7x – 9)

OpenStudy (texaschic101):

6(4x - 3) - 10x 24x - 18 - 10x 14x - 18 = ......for their to be no solution, you would have to make the x's cancel....and the integer cannot be -18. 14x - 18 = 2(7x - 10) 14x - 18 = 14x - 20 --- subtract 14x from both sides -18 = -20 ---- incorrect....no solutions 14x - 18 = 14x - 15 --- subtract 14x from both sides -18 = -15 --- incorrect....no solutions any questions ?

OpenStudy (anonymous):

Thank you! No, I got it :D

OpenStudy (texaschic101):

glad to help :)